OpenStudy (anonymous):

find the first four terms of this geometric sequence a1=3 and r=-3

OpenStudy (anonymous):

a , ar, ar^2, ar^3 a1 = a a2 = ar a3 = ar^2 a4 = ar^3

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

a1 = 3 a2 = a*r a2 = 3*(-3) = -9 a3 = 3*(-3)^2 = 3*9 = 27 a4 = 3*(-3)^3 = 3*(-27) = -81